OpenStudy (anonymous):

solve he equation x^3-x-1=0 show your method help!!!

OpenStudy (anonymous):

x^3-x-1 Taking x common : x(x^2-1)-1=0 Therefore : x=+-1

OpenStudy (mertsj):

It has no rational roots. 1 and -1 do not work.

OpenStudy (anonymous):

you mean by x=+-1????

OpenStudy (mertsj):

He means positive 1 and negative 1

OpenStudy (anonymous):

not a solution right?

OpenStudy (mertsj):

Right. Neither one of them make the sentence true.

OpenStudy (mertsj):

There are no rational roots.

OpenStudy (anonymous):

how u know that?

OpenStudy (mertsj):

The possible rational roots are the factors of the constant term divided by the factors of the coefficient of the leading term. Since both numbers are 1, the only possible rational roots are 1 and negative 1 and neither of those is a root.

OpenStudy (mertsj):

Do you know the rational roots theorem? e Rational Roots (or Rational Zeroes) Test is a handy way of obtaining a list of useful first guesses when you are trying to find the zeroes (roots) of a polynomial. Given a polynomial with integer (that is, positive and negative "whole-number") coefficients, the possible (or potential) zeroes are found by listing the factors of the constant (last) term over the factors of the leading coefficient, thus forming a list of fractions. This listing gives you a list of potential rational (fractional) roots to test -- hence the name of the Test.

OpenStudy (anonymous):

ok! any idea how to do it?

OpenStudy (mertsj):

I would test values to see where it crosses the x axis and then drill down.

OpenStudy (mertsj):

If x = 0, y is negative if x = 1, y is negative if x = 2, y is positive So there is a root between 1 and 2 if x = 1.5, y is positive so there is a root between 1 and 1.5

OpenStudy (anonymous):

it did actually for x between 1 and 1.5

OpenStudy (mertsj):

You could graph it with a graphing calculator and find the root also.

OpenStudy (anonymous):

with a graphing calculator they just give the solution approximately i do more care about the exact form of the solution and the method i need it to fin padovan's general term

OpenStudy (mertsj):

You could say: x^3-x=1 3st = -1 s=-1/3t s^3-t^3=1

OpenStudy (anonymous):

sorry i didnt understand what u wrote

OpenStudy (mertsj):

It is a method of changing the equation into a quadratic form with t^6 and t^3 so that it can be solved using the quadratic formula.

OpenStudy (mertsj):